# The Hilbert-space structure of free fermions in disguise

### v1

@@ IV.6 $M=6k+5$
Once again, we can construct a family $\{\tilde{Y}_{n}\}$ obeying (21). Together with (35) and (36), we thus have a total of $12\times(3k+1)+2=M-$ 1 operators, generating the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$.
@@
IV.7 Factoring out the symmetry: mapping to another FFD chain
@@
We illustrate this fact by projecting onto the eigenspaces of some of the symmetry generators, namely $\widetilde{h}_{3l+1}=X_{3l+1}Z_{3l+2}Z_{3l+3}$. For concreteness we choose to project onto the subspace where they all have eigenvalue $+1$ (a similar discussion holds for any choice of signs). Each block of three spins on sites $33l+1,3l+2,3l+$ 3 in the original representation can therefore be projected onto a macro-site of two spins, with basis states
@@
Here, 00 and $1$ correspond to the two eigenvalues of $Z_{3l+2}\,,Z_{3l+3}$. For a given choice of the eigenvalues, the state on site $13l+$ 1 is fixed to the eigenstate of $X_{3l+1}$ such that $X_{3l+1}Z_{3l+2}Z_{3l+3}=1$.
@@
[…]
Going further, we could “factor out” the rest of the symmetry algebra by fixing the eigenvalues of the remaining generators $\widetilde{h}_{6l+2}\widetilde{h}_{6l+3}$. The resulting model has a Hilbert space of dimension $\sim 2^{M/2}$ and we expect that it coincides with the two-chain model presented in Ref [19]. Since this direction is beyond the scope of this work, we will not pursue it further, but it remains an interesting question for future work.
V Ancillary free fermions in disguise
@@
V.1 The (super)symmetry algebra
@@
$\chi_{m}=a_{m}\leavevmode\nobreak\ \chi^{\prime}\prod_{i=1}^{m-1}h_{i}\,,\qquad m=1,\ldots M+1\,,$ (40)
where $a_{m}=\frac{b_{m}b_{m+2}\ldots}{b_{m+1}b_{m+3}\ldots}$, with the dots indicating products running until the end of the chain, that is $M$ or $M-1$ depending on the parity of $m$. The reason for the presence of $\chi^{\prime}$ is that $\chi_{m}$ defined as such commutes with all $\{h_{m}\}_{m=1,\ldots M}$ but two (namely $h_{m-2}$ and $h_{m+1}$), making it very easy to write its commutation relations with $H$, and therefore to find linear combinations of the $\chi_{m}$ that commute with it.
@@
Having defined the operators $O^{(n)}$, $E^{(n)}$, let us now study their commutation relations with the physical fermions. As is clear from the definition, all the $\chi_{m}$ commute with the right boundary mode $\chi$, except for the last, $\chi_{M+1}=\chi^{\prime}h_{M}$. As a result, out of the operators $O^{(n)}$, $E^{(n)}$, all those whose definition does not involve $\chi_{M+1}$ commute with $\chi$, and therefore with all fermion creation/annihilation operators. This corresponds to the $O^{(n)}$ for odd $M$, and to the $E^{(n)}$ for even $M$.
@@
V.2 Commuting transfer matrices
@@
$\mathcal{E}_{2m}(u)$ $=\mathcal{E}_{2m-2}(u)+u\leavevmode\nobreak\ \mathcal{O}_{2m-2}(u)\chi_{2m}\,,$ (46a)
$\mathcal{O}_{2m}(u)$ $=\mathcal{O}_{2m-2}(u)+u\leavevmode\nobreak\ \mathcal{E}_{2m-4}(u)\chi_{2m-1}\,,$ (46b)
$\mathcal{O}_{2m+1}(u)$ $=\mathcal{O}_{2m-1}(u)+u\leavevmode\nobreak\ \mathcal{E}_{2m-1}(u)\chi_{2m+1}\,,$ (46c)
$\mathcal{E}_{2m+1}(u)$ $=\mathcal{E}_{2m-1}(u)+u\leavevmode\nobreak\ \mathcal{O}_{2m-3}(u)\chi_{2m}\,.$ (46d)
@@ V.3 The ancillary fermions
which are the analog of the products $T(u)T(-u)$ in Fendley’s original construction. Those are polynomials in $u$, and obey the recursion relations
$\mathbb{E}_{2m}^{\pm}(u)$ $=\mathbb{E}_{2m-2}^{\pm}(u)+u\leavevmode\nobreak\ \mathbb{O}_{2m-2}^{\mp}(u)a_{2m}^{2}$ (50a)
$\mathbb{O}_{2m}^{\mp}(u)$ $=\mathbb{O}_{2m-2}^{\mp}(u)+u\leavevmode\nobreak\ \mathbb{E}_{2m-4}^{\pm}(u)a_{2m-1}^{2}$ (50b)
$\mathbb{O}_{2m+1}^{\pm}(u)$ $=\mathbb{O}_{2m-1}^{\pm}(u)+u\leavevmode\nobreak\ \mathbb{E}_{2m-1}^{\mp}(u)a_{2m+1}^{2}$ (50c)
$\mathbb{E}_{2m+1}^{\mp}(u)$ $=\mathbb{E}_{2m-1}^{\mp}(u)+u\leavevmode\nobreak\ \mathbb{O}_{2m-3}^{\pm}(u)a_{2m}^{2}\,.$ (50d)
@@
±$\pm\epsilon^{\prime}_{1}\pm\epsilon^{\prime}_{2}\ldots\pm\epsilon^{\prime}_{S^{\prime}}\,,$ (51)
@@
$\Psi^{\prime}_{\pm k}\propto\begin{cases}\mathcal{E}_{M}(\sqrt{\pm u_{k}})\leavevmode\nobreak\ E\leavevmode\nobreak\ \mathcal{E}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ even}\\ \mathcal{O}_{M}(\sqrt{\pm u_{k}})\leavevmode\nobreak\ O\leavevmode\nobreak\ \mathcal{O}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ odd}\end{cases}\,,$ (52)
@@
V.4 The symmetry algebra $\mathcal{A}_{F^{\prime}}$ in terms of the ancillary fermions
@@
We are therefore ready to state our final proposition, namely that the algebra $\mathcal{A}_{F^{\prime}}$ is generated by all bilinears in the ancillary fermions $\Psi^{\prime}_{k}$. This is confirmed by a counting of dimensions presented in Table 1, cf. Sec. III, where it is checked that the ancillary fermionic modes, together with the algebra $\mathcal{A}_{\widetilde{D}}$ of Sec. IV, indeed accounts for all the degeneracies observed in the spectrum.
VI Outlook

### v2

@@ 4.6 $M=6k+5$
Once again, we can construct a family $\{\tilde{Y}_{n}\}$ obeying (21). Together with (35) and (36), we thus have a total of $2\times(3k+1)+2=M-1$ operators, generating the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$.
@@
4.7 Factoring out the symmetry: mapping to another FFD chain
@@
We illustrate this fact by projecting onto the eigenspaces of some of the symmetry generators, namely $\widetilde{h}_{3l+1}=X_{3l+1}Z_{3l+2}Z_{3l+3}$. For concreteness we choose to project onto the subspace where they all have eigenvalue +1+1 (a similar discussion holds for any choice of signs). Each block of three spins on sites $3l+1,3l+2,3l+3$ in the original representation can therefore be projected onto a macro-site of two spins, with basis states
@@
Here, 0 and $1$ correspond to the two eigenvalues of $Z_{3l+2}\,,Z_{3l+3}$. For a given choice of the eigenvalues, the state on site $3l+1$ is fixed to the eigenstate of $X_{3l+1}$ such that $X_{3l+1}Z_{3l+2}Z_{3l+3}=1$.
@@
where on each macro-site $X,Z$, and $X^{\prime},Z^{\prime}$ refer respectively to the action on the first or second spin. Depending on the value of $M({\rm mod}~3)$, the mapping (39c) may need to be adapted for the rightmost operators $h_{m}$, but we will not discuss this in detail here. In summary, the projection has allowed us to map the model onto a chain of $\sim 2M/3$ spins, with the same energy spectrum as the original model but less degeneracies.
Going further, we could “factor out” the rest of the symmetry algebra by fixing the eigenvalues of the remaining generators $\widetilde{h}_{6l+2}\widetilde{h}_{6l+3}$. The resulting model has a Hilbert space of dimension $\sim 2^{M/2}$ and we expect that it coincides with the two-chain model presented in Ref [19] and worked out in more detail in Ref.  [27] (in that reference, “minimal representations" of the FFD algebra were presented for any value of $M$). Since this direction is beyond the scope of this work, we will not pursue it further, but it remains an interesting question for future work.
5 Ancillary free fermions in disguise
@@
5.1 The (super)symmetry algebra
@@
$\chi_{m}=a_{m}~\chi^{\prime}\prod_{i=1}^{m-1}h_{i}\,,\qquad m=1,\ldots M+1\,$, (40)
where $a_{m}=\frac{b_{m}b_{m+2}\ldots}{b_{m+1}b_{m+3}\ldots}$, with the dots indicating products running until the end of the chain, that is $M$ or M1M-1 depending on the parity of $m$. The reason for the presence of $\chi^{\prime}$ is that $\chi_{m}$ defined as such commutes with all $\{h_{m}\}_{m=1,\ldots M}$ but two (namely $h_{m-2}$ and $h_{m+1}$), making it very easy to write its commutation relations with $H$, and therefore to find linear combinations of the $\chi_{m}$ that commute with it.
@@
Having defined the operators $O^{(n)}$, $E^{(n)}$, let us now study their commutation relations with the physical fermions. As is clear from the definition, all the $\chi_{m}$ commute with the right boundary mode $\chi$, except for the last, $\chi_{M+1}=\chi^{\prime}h_{1}\ldots h_{M}$. As a result, out of the operators $O^{(n)}$, $E^{(n)}$, all those whose definition does not involve $\chi_{M+1}$ commute with $\chi$, and therefore with all fermion creation/annihilation operators. This corresponds to the $O^{(n)}$ for odd $M$, and to the $E^{(n)}$ for even $M$.
@@
5.2 Commuting transfer matrices
@@
$\mathcal{E}_{2m}(u)$ $=\mathcal{E}_{2m-2}(u)+u~\mathcal{O}_{2m-2}(u)\chi_{2m}\,$, (46a)
$\mathcal{O}_{2m}(u)$ $=\mathcal{O}_{2m-2}(u)+u~\mathcal{E}_{2m-4}(u)\chi_{2m-1}\,$, (46b)
$\mathcal{O}_{2m+1}(u)$ $=\mathcal{O}_{2m-1}(u)+u~\mathcal{E}_{2m-1}(u)\chi_{2m+1}\,$, (46c)
$\mathcal{E}_{2m+1}(u)$ $=\mathcal{E}_{2m-1}(u)+u~\mathcal{O}_{2m-3}(u)\chi_{2m}\,.$ (46d)
@@ 5.3 The ancillary fermions
which are the analog of the products T(u)T($u)T($-u) in Fendley’s original construction. Those are polynomials in $u$, and obey the recursion relations
$\mathbb{E}_{2m}^{\pm}(u)$ $=\mathbb{E}_{2m-2}^{\pm}(u)+u~\mathbb{O}_{2m-2}^{\mp}(u)a_{2m}^{2}$ (50a)
$\mathbb{O}_{2m}^{\mp}(u)$ $=\mathbb{O}_{2m-2}^{\mp}(u)+u~\mathbb{E}_{2m-4}^{\pm}(u)a_{2m-1}^{2}$ (50b)
$\mathbb{O}_{2m+1}^{\pm}(u)$ $=\mathbb{O}_{2m-1}^{\pm}(u)+u~\mathbb{E}_{2m-1}^{\mp}(u)a_{2m+1}^{2}$ (50c)
$\mathbb{E}_{2m+1}^{\mp}(u)$ $=\mathbb{E}_{2m-1}^{\mp}(u)+u~\mathbb{O}_{2m-3}^{\pm}(u)a_{2m}^{2}\,.$ (50d)
@@
$\pm\epsilon^{\prime}_{1}\pm\epsilon^{\prime}_{2}\ldots\pm\epsilon^{\prime}_{S^{\prime}}\,$, (51)
@@
$\Psi^{\prime}_{\pm k}\propto\begin{cases}\mathcal{E}_{M}(\sqrt{\pm u_{k}})~E~\mathcal{E}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ even}\\ \mathcal{O}_{M}(\sqrt{\pm u_{k}})~O~\mathcal{O}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ odd}\end{cases}\,$, (52)
@@
5.4 The symmetry algebra $\mathcal{A}_{F^{\prime}}$ in terms of the ancillary fermions
@@
We are therefore ready to state our final proposition, namely that the algebra $\mathcal{A}_{F^{\prime}}$ is generated by all bilinears in the ancillary fermions $\Psi^{\prime}_{k}$. This is confirmed by a counting of dimensions presented in Table 1, cf. Sec. 3, where it is checked that the ancillary fermionic modes, together with the algebra $\mathcal{A}_{\widetilde{D}}$ of Sec. 4, indeed accounts for all the degeneracies observed in the spectrum.
5.5 The remaining zero mode for $M=6k+3$
As mentioned above, for the special case $M=6k+3$, the algebras $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$ are enough to lift all degeneracy factors of the spectrum, but one. This can be explained by the existence of an additional zero mode, which commutes both with the physical fermions and with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$. For $M=3$, such an operator can easily be found as $(h_{1}+h_{2})\chi_{M}$. For $M=9$, it can be found as a bilinear in the $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. For general $M=6k+3$ we conjecture that an operator of such a form should exist, namely some multilinear combination of $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. Furthermore, it should be possible to relate the existence of this operator to the zero modes evidenced in [29] more generally for all $M\notin 3\mathbb{N}+1$. However, a difference between our approach and the zero modes of [29] is that the latter do not necessarily commute with the rest of the symmetry algebra, that is, with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$.
6 Outlook